OpenStudy (anonymous):

find the values of a,b and c that make the function \[f(r) = ar^{2} + br + c; \] \[r<0 , 2\sin^{2}(r) + 3\cos^2(r)\] \[r \ge 0\] twice differentiable at r=0.

OpenStudy (anonymous):

hard question... have no idea about how to solve it

OpenStudy (anonymous):

what is the second function?

OpenStudy (anonymous):

thats what is given in the question :D

OpenStudy (anonymous):

err its ODTU universities calculus 1 final exam question

OpenStudy (anonymous):

do they represent value of r

OpenStudy (anonymous):

satellite, let me repost question?

OpenStudy (anonymous):

well a first though is that if it is going to be differentiable at 0 at least it better be continuous there. and further since \[f'_-(r)=2ar+b\] and \[f'_-(0)=0\] it should also be the case that \[f'_+(0)=4\sin(r)\cos(r)-6\cos(r)\sin(r)=0\]

OpenStudy (anonymous):

so i am thinking that the first thing we know is that \[c=3\]

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

it is not obvious to me how we get a and b because both derivatives are 0 at 0

OpenStudy (anonymous):

hmm i get this question from ODTU's final calculus exam

OpenStudy (anonymous):

maybe from the second the second derivative? lets see

OpenStudy (anonymous):

mybe, lets try

OpenStudy (anonymous):

lets call some other people for help? the more brains focus in one question, the easier we get the result

OpenStudy (anonymous):

i wrote this out in another window! let me repost

OpenStudy (anonymous):

hi korcan, lets see from f '(r)=4sin r cos r -6cos r sinr =-2sin r cos r=0, r=720 degrees and f ''(r)=-2[-sin r sin r+cos r cos r] =2[(cos r)^2 -(sin)^2]=0, r=45 degrees

OpenStudy (anonymous):

lets build up the puzzle little by little...lol

OpenStudy (anonymous):

err mark o. we satellite already solved it :D (reposted the question )

OpenStudy (anonymous):

can you have a look at my other question? its at left :D starts with "given that..."

OpenStudy (anonymous):

ah ok ,,you did? where at